OpenStudy (anonymous):

i(x)=1−3×10^(2x+3) I have the range and domain as being x all the reals and y less than one but I don't understand why I have started doing the inverse of the function but I am stuck and the person who was helping me has gone offline :(

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you don't know why \[1-3\times 10^{2x+3}<1\]?

OpenStudy (misty1212):

\[10^{2x+3}\] is positive, so \[-3\times 10^{2x+3}\] is negative

OpenStudy (misty1212):

that makes the whole thing always less that one

OpenStudy (anonymous):

ahhhh ok :)

OpenStudy (anonymous):

Are you able to help with the inverse?

OpenStudy (misty1212):

sure why not?

OpenStudy (misty1212):

\[1-3\times 10^{2x+3}=y\\ -3\times 10^{2x+3}=y-1\\ 10^{2x+3}=\frac{1-y}{3}\\ 2x+3=\log(\frac{1-y}{3})\]

OpenStudy (misty1212):

then you solve for \(x\) \[2x=\log(\frac{1-y}{3})-3\\ x=\frac{1}{2}\log(\frac{1-y}{3})-\frac{3}{2}\]

OpenStudy (anonymous):

Awwww thank you so very much

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

can I ask does this function have asymptotes?